# Global dynamical structures from infinitesimal data

### v1

@@ Abstract
Discovering mechanisms underlying the behaviors of complex, high dimensional, and nonlinear dynamical systems is a central goal of the natural and synthetic sciences. Breakthroughs in machine learning in concert with increasing capacities for computation and data collection have enabled the use of trajectory measurements for learning predictive models. However, rigorous approaches for interpreting mechanisms from such models remain elusive, and asymptotic prediction accuracy suffers if the model does not capture important state space structures (e.g., attracting invariant sets). These limitations are especially pressing for system-level behaviors such as whole-body locomotion, where discontinuous, transient, and multiscale phenomena are common and prior models are rare. To take the next step towards a theory and practice for dynamical inference of complex multiscale systems in biology and beyond, we introduce VERT, a framework for learning the attracting sets that characterize global system behavior without recourse to learning a global model. Our approach is based on an infinitesimal-local-global (ILG) framework for estimating the proximity of any sampled state to the attracting set, if one exists, with formal accuracy guarantees. We demonstrate our approach on phenomenological and physical oscillators with hierarchical and impulsive dynamics, finding sensitivity to both global and intermediate attractors composed in sequence and parallel. Application of VERT to human running kinematics data reveals insight into control modules that stabilize task-level dynamics, supporting a longstanding neuromechanical control hypothesis. The VERT framework thus enables rigorous inference of underlying dynamical structure even for systems where learning a global dynamics model is impractical or impossible.
@@
Recent breakthroughs in applied machine learning have seen the development of sophisticated approaches that leverage measured trajectory data, conceived as the output of some hidden dynamical system, and using those observations to learn a model that minimizes a system-specific prediction error. Linear [2, 3] and nonlinear [4, 5, 6, 7] dimensionality reduction can reveal latent variables that capture variance, but do not explicitly account for the temporal structure of trajectory datasets that contains critical information about underlying dynamics. Both can fail to capture structure necessary for understanding control mechanisms [8]. Another paradigm conjectures that observed trajectories are the outputs of dynamical systems and uses trajectory observations to learn a model that minimizes a system-specific prediction error. Exemplar approaches employ a range of approximation classes including linear dynamic modes, sets of nonlinear basis functions, and neural networks, and show that impressive prediction accuracy can be achieved by formulating the learning problem using insight from methods including Koopman theory [9, 10], symbolic regression [11, 12], and sparse regularization [12]. Further modifications to the model fitting procedure can improve accuracy by incorporating inductive priors from dynamical systems theory. For example, the loss function used for model training can be modified to promote more parsimonious coordinate representations [13], or to improve model reduction in systems with multiple time scales when a full order model is available [14]. A recent foundational approach uses spectral submanifold theory to learn reduced models of nonlinear dynamics as extended normal forms, demonstrating significant improvements to prediction accuracy for mechanical oscillations [15]. However, if the observations are not well approximated by a known (or hypothesized) continuous dynamical system, the need to select a global approximation class and learning criteria a priori presents a significant obstacle to learning an interpretable model that faithfully reproduces underlying mechanisms. The limitations of global model identification can be reduced by temporal segmentation of the trajectories into windows that are well approximated by linear dynamical systems [16, 17], or spatial segmentation of the domain into local charts that are more amenable to approximations of the nonlinear dynamics [18]. However, transitions between dynamical modes can reflect important mechanisms such as changing of constraints or internal control processes rather than limitations of the modeling apparatus [^1] For example, the making and breaking of contact by appendages during animal and robot locomotion induces transitions between characteristically nonlinear dynamical regimes, motivating the necessity of considering infinitesimal stability properties in determining where qualitative changes in behavior occur..
@@
Leveraging insights from differential topology [20] applied to the structure of hyperbolic invariant manifolds [21], we introduce the Vielbein Recovery of Templates (VERT) framework, an infinitesimal-local-global (ILG) paradigm for localizing persistent attracting submanifolds in output (observation) coordinates (Fig. 1). The infinitesimal component manifests itself in the assumption on the jets of the underlying dynamics, stipulating a decomposition of the tangent space into (integrable) distributions of contracting and neutral subspaces, which can be inferred from local data. The local existence of the integral submanifolds is phenomenologically strengthened to the existence of the global low-dimensional attractors and attendant basins of attractions [22]  [^2] We remark that the VERT framework is applicable in hyperbolic dynamics where local contracting/expanding with respect to the offered Riemannian norm occurs only in some parts of the attractor (they need to occur somewhere, however, to satisfy the global hyperbolicity assumptions [55]). Once detected there, the normal hyperbolicity in broader regions can be inferred by extension along the trajectories. As an example, we cite van der Pol oscillators (whose entire limit sets satisfy the global criteria for hyperbolicity of [21], but lose a local attraction property in the standard Riemannian norm in the regions where the energy is pumped into the system by the “Coriolis term”).. This broad framework shares with much of the prior literature the widely held premise that dynamical behavior of physical interest is typically characterized by a collapse of dimension – the presence of a low-dimensional attracting invariant submanifold onto which a system’s trajectories quickly converge within the nominally high dimensional physical state space.
@@
Motivated by this insight, our central contribution is the development of a theory and computational procedure for isolating from observed trajectory data the global locus of the lower dimensional attracting submanifolds that carry the long term system behavior. Critically for biological applications, our approach achieves this analysis while circumventing the need to fit a global (or piecewise global) model of the vector field to the trajectory data. Specifically, we show that a ’fiberwise distance estimator’ whose sublevel sets indicate trajectory segments on or close to an attracting set can be constructed from infinitesimal-scale, linear dynamics estimated from local samples in a pointwise manner, affording localization of these loci of characteristic behavior without prior commitment to any parameterized model or structural assumptions beyond normal hyperbolicity — a necessary condition for their persistence under perturbations. Using models of both continuous and hybrid dynamical systems, we demonstrate the efficacy of our approach for discovering structurally stable attracting submanifolds with a range of geometric and dynamical properties without recourse to any other prior assumptions about the nature of the full system or attractor. Finally, we illustrate the applicability of our framework to postural trajectory data from motion capture of human treadmill running whose future systematic pursuit promises to reveal insight into the neuromechanical control of the hybrid, nonlinear dynamics underlying sensorimotor behavior.
@@
Robust Behavior and Hyperbolic Attractors
@@
Figure 1: Overview of the VERT computational framework. Panels (A-B) depict the relation to motivating hierarchical control hypotheses for complex, multi-scale systems, viewed through an illustrative application setting of the postural dynamics of human running. Panels (D-G) outline the steps of the VERT algorithm. A. Stable task-level system behaviors such as running emerge from the coupled, highly nonlinear, closed loop dynamics of many body segments and their interaction with the environment. B. We hypothesize that for many tasks, the global task-level behavior may be decomposed into hierarchically arranged control modules, whose coordinated interactions characterize the dynamics of appendages and body segments. The control module state spaces are embedded as submanifolds (postures) that carry the constituent dynamics (templates) of the full, coordinated system. C. We can use VERT to learn a spatiotemporal filter on the trajectory data that isolates locally stable (template) behavior from transient behaviors reflecting disturbances and stabilizing control (anchoring). D. VERT estimates the distance from each sample to the attractor guiding its asymptotic behavior ($f(z)$, Def. 1). The estimation procedure for a point $z^{*}$ begins with identification of its metric nearest neighbors in state space. E. A Cartesian basis for the tangent space at $z^{*}$ of the manifold containing the trajectories is estimated. The local subsample of the vector field (temporal differences, gray arrows) is projected onto the $n$-dimensional local vielbein basis (black arrows). F. Using this representation, we can estimate the component of the vector field on the vertical subspace, the local trivialization of the global, typically nonlinear vertical subbundle of the attractor. We can use this estimate to approximate the distance from $z^{*}$ to its base point on the unknown attractor $\mathcal{A}$. G. Steps D-G are repeated for an appropriate number of samples. Filtering the trajectory data for sublevel sets of $f(z)$ reveals the posture carrying the template dynamics (blue) and hierarchical stabilization (anchoring) intervals of what subsequent hypothesis-informed analysis can posit as the stance leg and upper body posture during stance (See Fig. 5).
A common modeling assumption is that characteristic system behaviors lie on or near a lower dimensional, often nonlinear space of latent variables in the full observable state space. However, the presence of external disturbances and internal control processes can perturb the state away from the lower dimensional latent space. Robustness to such disturbances requires that the latent space is an attracting invariant set, i.e., perturbed trajectories return asymptotically. Motivated by applications to experimentally measured data, we assume that the attracting invariant sets are locally structurally stable, that is, persistent under noise and perturbation. Normal hyperbolicity is a sufficient condition for local structural stability; for modeling purposes, we assume that detectable attracting sets are normally hyperbolic invariant manifolds (NHIM) [21].
@@
Dynamic locomotion [24] provides a tractable behavior for exploring principles underlying robust, multiscale dynamical systems. The target behaviors of locomoting systems, typically translation and rotation of the body, are both observable and interpretable. A longstanding question in the sensorimotor control of locomotion is how the abundance of muscles and joints in the body are coordinated to achieve a large and diverse set of typically lower dimensional tasks. The pioneering work of Nikolai Bernstein [25] suggests that whole body movements are constructed generatively from modular ’submovements’. Modern electromyography (EMG) studies in both humans and non-human animals support the hypothesis that the neuromuscular system is organized into modular units (control modules) that can be recruited in series and parallel to accomplish high level motor tasks, and even recombined and fused through rehabilitative training when partial motor function is lost [26]. Advances in automated pose estimation provide the possibility of collecting large, rich kinematic trajectory datasets, affording the possibility of building surrogate data-driven models that reproduce these critical features of robust sensorimotor behavior.
The templates and anchors framework (TAF) [1] articulates the control module hypothesis in the context of biomechanics and dynamical systems theory, stipulating that the characteristic postures of locomotion tasks lie on attracting invariant submanifolds (posture submanifolds) of the full postural state space. The posture submanifold and its restriction dynamics reveal the template, a low dimensional dynamical system that captures the target behavior for the task. This low dimensional system is anchored in the full (postural) state space by stabilizing control forces that render the posture submanifold attracting and invariant. Beyond the setting of locomotion, constructing a complex dynamical system from simpler subsystems is a ubiquitous approach for system synthesis and analysis [27]. In the following sections, we review the general setting of dynamical systems constructed from parallel or sequential composition of subsystems, with motivating examples from dynamic locomotion and the TAF.
Parallel Composition of Attractors
In control module compositions from robotics, there is typically a great deal of choice in the geometry of the postures (the manner in which the template dynamics are embedded) and the relative strengths of the various anchoring gains that determine their relative time constants of attraction — and, thereby, the characteristic order in which trajectories approach the increasingly restricted succession of intersecting postures. Thus, from the viewpoint of robotics there is significant interest in not simply testing the TAF rendering of the control module composition hypothesis but in actually discovering what posture principles are associated with what kinds of compositions. This is the key question that our computational pipeline addresses.
To test the efficacy of our pipeline relative to this circle of ideas, we introduce a model of coupled Hopf-type oscillators anchored by a stable flow in a higher dimensional Euclidean space, the Anchored Hopf Model (A-HM). The structure of the coupled limit cycle modules of the A-HM is a first order phase oscillator, whose illustrated value is motivated by empirical and modeling studies of sprawled posture legged locomotion in insects [28], while the variously embedded postures (the coupled attracting invariant submanifolds) are motivated by the robotics examples [29, 30, 31, 32, 33]. Specifically, embedding of the coupled limit cycle system in the higher dimensional ’posture’ space by a stabilizing vector field is motivated by a recent empirically validated algorithm for anchoring planar behaviors in spatial legged robots [33]. For our computational test system (Eqns. 1 & 2), the gains of the control forces that drive the state towards the limit cycle and entrain the two phase oscillators are set by the scalar parameters $\gamma_{\rho}>0$ and $\gamma_{\theta}>0$, respectively.
The template dynamics we consider within a Euclidean state space (with state $x\in\mathbb{R}^{4}$) is easiest to render in polar coordinates:
$\begin{split}&\dot{\rho}_{i}=\gamma_{\rho}(1-\rho_{i})\rho_{i}\\ &\dot{\theta}_{i}=\omega+\gamma_{\theta}\sin(\theta_{i}-\theta_{j})\end{split}$ (1)
where $i,j\in\{1,2\}$. Observe that the dynamics of the phase variables decouples from the amplitudes $\rho$, and has the structure of the Kuramoto oscillator system. The system entrains to a global limit cycle so long as $\gamma_{\theta}>0$.
The reduced state space in which the limit cycles are anchored is in turn anchored in $\mathbb{R}^{D}$. For modeling purposes, we assume that the dynamics of the remaining coordinates is globally asymptotically stable e.g., is given by
$\dot{y}=Ay$ (2)
with a Hurwitz $A$, and $y=(y_{5},\ldots,y_{D})$ is the vector of the auxiliary variables. The state vector for the full system composed of Eqs. 1 & 2 is thus $(z_{1},...,z_{D})=(x_{1},...,x_{4},y_{5},...,y_{D})\in\mathbb{R}^{D}$. The rate of convergence to the anchored plane at $y=0$ is set by the real part of the eigenvalues $\sigma$ of $A$, $\Re\sigma=-\gamma_{\delta}$.
The behaviors realized by particular trajectories of the total system are determined by the relative magnitudes of the forces imparted by the control modules, the initial state, and the nature of perturbations to which the system is subjected, as in Fig. 2. The total system contains a hierarchy of attracting submanifolds whose appearance is governed by the relative strength of the gain parameters summarized in Table 1 that we now discuss. The lowest level of the hierarchy, towards which all initial conditions converge asymptotically, is a Hopf limit cycle, in the coupled case $\gamma_{\theta}>0$. (If the oscillators are decoupled, i.e., if $\gamma_{\theta}=0$, then the lowest dimensional attracting submanifold is the torus formed as the product space of the independent limit cycles.
This limit torus is anchored in the 4D subspace of the Hopf systems by the subsystems in Eq. 2, if the spectrum of $A$ lies in the half-plane of $\{\Re\sigma<-\gamma_{\rho}\}$. If this is not the case, the state may converge to an annular subdomain of $\mathbb{R}^{D}$ before reaching the 4D Hopf subspace, providing multiple qualitatively distinct routes for anchoring the lowest level template in the full system.
Figure 2: Visualization of parallel anchoring hypotheses for the A-HM system described by Eqs. 1 & 2, with $\gamma_{\theta}\gg 0$. From the initial state $z(0)$, each trajectory approaches the limit cycle $\mathcal{S}$ (red) asymptotically, represented here by $z(t_{S})$. The relative values of control parameters ($\gamma_{\rho},\gamma_{\delta}$) set the sequence of hierarchically arranged intermediate attractors that determine the behavior of the pre-limit cycle trajectory. Following Table 1: $\mathbf{H_{0}}$ in red, $\mathbf{H_{1}}$ in cyan, $\mathbf{H_{2}}$ in purple, and $\mathbf{H_{3}}$ in black. Trajectory segments on their respective intermediate attractors are highlighted in blue (c.f., Table 1, Column 2).
Subsystem Gain Anchoring Hierarchy Hypothesis VERT Detection Hypothesis
$(\mathbf{H_{0}})$ $\gamma_{\theta}\gg 0$ $z_{i}\in\mathbb{R}^{d}\rightarrow\mathcal{S}\ni z_{f}$ Detect $\mathcal{S}$
Table 1: Table of hypothesized anchoring mechanisms ($\mathbf{H_{0-3}}$) for the Hopf system. We hypothesize the sequence of hierarchically embedded submanifolds that typical trajectories approach for each system configuration, and the expected behavior of the VERT pipeline if the hierarchy is detected.
Sequential Composition of Attractors
Analogous to the composition of control modules in state space described in the previous section, changes in target behavior or environment often necessitate the sequential composition of control modules in time. Achieving goal states while avoiding obstacles by back-chaining basins of attraction around successively targeted attractors by either design choice [34] or as necessitated by limited control affordance [35] represents a common instance of such nonsmooth, event-driven switches between stable continuous dynamics in the engineering control literature. However, even steady state locomotion behaviors, e.g., running with legs, involve impulsive changes in dynamics over the course of a single period as limbs make and break contact with the substrate. Each transition incurs the possibility of disturbances and postural errors that must be stabilized either by passive dynamics or active feedback control. The theory of hybrid dynamical systems provides a rigorous framework for the analysis and control of a broad array of such physically common systems that are parsimoniously modeled by imposing nonsmooth resets guarded by event-driven switches between distinct, smooth dynamical regimes [36, 37].
Relaxation oscillators offer a parametrically graded family of limit cycles that facilitate the study of nonlinear dynamics as they approach and begin to exhibit hybrid dynamical behavior [38]. In such systems, trajectories on the attracting limit cycle alternate between slow convergence along the cycle towards a near-equilibrium condition (relaxation) and fast, impulsive switches in the position of the equilibrium [38]. The van der Pol system is a canonical example of a relaxation oscillator when its bifurcation parameter that quantifies the strength of nonlinear damping exceeds a threshold. The alternation between slow, locally hyperbolic attracting dynamics and impulsive switches in the location of the local attractor provide an interpretable and relatively simple model in which to explore the anchoring of sequentially composed templates.
We use the anchored van der Pol (A-vdP) system, the cross product vector field of second order vector fields corresponding to the van der Pol system and an overdamped harmonic oscillator, as a testing example. In Cartesian coordinates:
$\begin{split}&\dot{x}_{1}=x_{2}\\ &\dot{x}_{2}=-\sin(x_{1})+\mu(\cos^{2}(x_{1})-0.5)x_{2}\\ &\dot{x}_{3}=x_{4}\\ &\dot{x}_{4}=-x_{3}-bx_{4}\end{split}$ (3)
where the plane with coordinates $(x_{1},x_{2})$ contains the van der Pol limit cycle, and $(\dot{x}_{3},\dot{x}_{4})$ defines a (decoupled) vector field on $(x_{3},x_{4})$ that anchors the limit cycle in $\mathbb{R}^{4}$ when the linear damping term $b$ is positive. When the nonlinear damping $\mu$ is sufficiently large, the system exhibits relaxation oscillations.
@@
We introduce the Vielbein Recovery of Templates (VERT) framework, an ILG approach for identifying hierarchically embedded attracting sets from trajectory data. VERT takes a set of observed trajectories and returns the proximity of each point to the normally attracting invariant manifold (NAIM) that governs its asymptotic behavior. To establish a theoretical foundation for the computational heuristics used by VERT, we model the trajectories $\{z_{i}(t),t\in[0,T_{i}]\}^{N}_{i=1}$ as samples from the (possibly noisy) flow of a dynamical system $(\mathcal{M},\xi)$, where $\mathcal{M}$ is a piecewise smooth Riemannian manifold of dimension $n$, and $v:\mathcal{M}\rightarrow T\mathcal{M}$ is a piecewise $C^{r\geq 1}$ vector field. We assume the trajectories are embedded in a Euclidean observation space $\mathbb{R}^{D}$.
@@
Let $\mathcal{A}\subset\mathcal{M}$ be an invariant submanifold of dimension $a$ for $(\mathcal{M},v)$: $\phi_{t}(\mathcal{A})\subset\mathcal{A},t>0$, for $\phi_{\cdot}$ being the flow associated to the vector field $v$. We will denote by $U\supset\mathcal{A}$ a tubular neighborhood of this submanifold [20] which is attracted to $\mathcal{A}$. Normal Hyperbolicity implies that the trajectories starting in the neighborhood $U$ converge to $\mathcal{A}$, exponentially fast, uniformly in some fixed Riemannian structure. We can introduce, locally, a privileged coordinate system $z=(x,y)$ on $U$, such that the the coordinate $x$ parameterizes the points of $\mathcal{A}$, while the attractor is given by $\mathcal{A}=\{y=0\}$. The fibers $\mathcal{F}_{x}$ of the projection $(x,y)\mapsto x$ are the analogues of the stable manifolds for the hyperbolic dynamics, trivialized through the selected coordinate system. We notice that this coordinate system does not have to (and in fact, almost never does) extend to a global one. However, its usage provides a convenient modeling setting, and leads to the heuristic assumptions outlined below  [^3] We use the German term Vielbein for frame, a local coordinate system to avoid contamination with the video processing context: the term is used in this meaning in the physics literature [56]..
@@
$v_{x}=\begin{pmatrix}u_{0}+S(x,y)\\ A_{1}y+T(x,y)\end{pmatrix}=\begin{pmatrix}u_{0}\\ A_{1}y\end{pmatrix}+B(x,y).$ (4)

### v2

# Global dynamical structure from infinitesimal data

@@ Abstract
Scientists and engineers alike target modeling of complex, high dimensional, and nonlinear dynamical systems as a central goal. Machine learning breakthroughs alongside mounting computation and data advance the efficacy of learning from trajectory measurements. However scientifically interpreting data-driven models – e.g., localizing attracting sets and their basins – remains elusive. Such limitations particularly afflict identification of system-level regulatory mechanisms characteristic of living systems – e.g., stabilizing control for whole-body locomotion – where discontinuous, transient, and multiscale phenomena are common and prior models are rare. As a next step towards theory-grounded discovery of behavioral mechanisms in biology and beyond, we introduce VERT, a framework for discovering attracting sets from trajectories without recourse to any global model. Our infinitesimal-local-global (ILG) pipeline estimates the proximity of any sampled state to an attracting set, if one exists, with formal accuracy guarantees. We demonstrate our approach on phenomenological and physical oscillators with hierarchical and impulsive dynamics, finding sensitivity to both global and intermediate attractors composed in sequence and parallel. Application of VERT to human running kinematics data reveals insight into control modules that stabilize task-level dynamics, supporting a longstanding neuromechanical control hypothesis. The VERT framework promotes rigorous inference of underlying dynamical structure even for systems where learning a global dynamics model is impractical or impossible.
preprint: APS/123-QED
I Introduction
@@
Linear [52, 10] and nonlinear [2, 17, 14, 38] dimensionality reduction can reveal spatial patterns and latent variables that capture variance, but do not explicitly account for the temporal structure of trajectory datasets and are thus insufficient for assessing, e.g., topological invariants of the underlying flow and their stability. Both can fail to capture structure necessary for understanding control mechanisms [60]. Recent breakthroughs in applied machine learning have seen the development of sophisticated approaches that conceive trajectories as the output of a hidden dynamical system, and then fit a model that minimizes a system-specific prediction error. Approaches that parameterize the dynamics as linear modes, sets of nonlinear basis functions, or deep neural networks have demonstrated impressive prediction accuracy by incorporating insights from Koopman theory [59, 41], symbolic regression [47, 7], and sparse regularization [7, 9]. Spectral submanifold theory can be used to learn reduced models of nonlinear dynamics as extended normal forms, demonstrating significant improvements to prediction accuracy for mechanical oscillations [8].
This synthesis of machine learning and dynamical system theory has demonstrated promise for enabling discoveries in physical systems. However, data-driven insight into the transient-rich and non-stationary dynamics typical of biological systems has remained relatively elusive. Inductive biases from invariant manifold theory have enabled accurate learning of reduced models in multiscale systems with transients when a full order model is available in addition to observations [42]. Heuristics can be used to segment trajectories or the model domain into temporal windows that are well-approximated by linear systems [15, 11] or spatial subdomains that are amenable to nonlinear approximations [20]. However, transitions between dynamical modes can reflect important mechanisms such as changing of constraints or internal control processes rather than limitations of the modeling apparatus [^1] For example, the making and breaking of contact by appendages during animal and robot locomotion induces transitions between characteristically nonlinear dynamical regimes, motivating the necessity of considering infinitesimal stability properties in determining where qualitative changes in behavior occur.. Principled tools for inferring dynamical structure from trajectory data without recourse to fitting a global model would extend the range of settings where explicit model learning could be applied, in addition to providing quantitative insights into systems where learning an explicit dynamics model is impractical.
@@
Leveraging insights from differential topology [27] applied to the structure of hyperbolic invariant manifolds [26], we introduce the Vielbein Recovery of Templates (VERT) framework, an ILG paradigm for localizing persistent attracting submanifolds in output (observation) coordinates (Fig. 1). The infinitesimal component manifests itself in the assumption on the jets of the underlying dynamics, stipulating a decomposition of the tangent space into (integrable) distributions of contracting and neutral subspaces, which can be inferred from local data. The local existence of the integral submanifolds is phenomenologically strengthened to the existence of the global low-dimensional attractors and attendant basins of attraction [23]  [^2] We remark that the VERT framework is applicable in hyperbolic dynamics where local contracting/expanding with respect to the offered Riemannian norm occurs only in some parts of the attractor (they need to occur somewhere, however, to satisfy the global hyperbolicity assumptions [51]). Once detected there, the normal hyperbolicity in broader regions can be inferred by extension along the trajectories. As an example, we cite van der Pol oscillators (whose entire limit sets satisfy the global criteria for hyperbolicity of [26], but lose a local attraction property in the standard Riemannian norm in the regions where the energy is pumped into the system by the “Coriolis term”).. This broad framework shares with much of the prior literature the widely held premise that dynamical behavior of physical interest is typically characterized by a collapse of dimension – the presence of a low-dimensional attracting invariant submanifold onto which a system’s trajectories quickly converge within the nominally high dimensional physical state space.
@@
Motivated by this insight, our central contribution is the development of a theory and computational procedure for isolating from observed trajectory data the global locus of the lower dimensional attracting submanifolds that carry the long term system behavior. Critically for biological applications, our approach achieves this analysis while circumventing the need to fit a global (or piecewise global) model of the vector field to the trajectory data. Specifically, we show that a ’fiberwise distance estimator’ whose near-vanishing sublevel sets indicate trajectory segments on or close to an attracting set can be constructed from infinitesimal-scale, linear dynamics estimated from local samples in a pointwise manner, affording localization of these loci of characteristic behavior without prior commitment to any parameterized model or structural assumptions beyond normal hyperbolicity — a necessary condition for their persistence under perturbations. Using models of both continuous and hybrid dynamical systems, we demonstrate the efficacy of our approach for discovering structurally stable attracting submanifolds with a range of geometric and dynamical properties without recourse to any other prior assumptions about the nature of the full system or attractor. Finally, we illustrate the applicability of our framework to postural trajectory data from motion capture of human treadmill running whose future systematic pursuit promises to reveal insight into the neuromechanical control of the hybrid, nonlinear dynamics underlying sensorimotor behavior.
@@
State Space Structures of Multiscale Dynamical Systems
@@
Figure 1: Overview of the VERT computational framework. Panels (A-B) depict the relation to motivating hierarchical control hypotheses for complex, multi-scale systems, viewed through an illustrative application setting of the postural dynamics of human running. Panels (D-F) outline the steps of the VERT algorithm. A. Stable task-level system behaviors such as running emerge from the coupled, highly nonlinear, closed loop dynamics of many body segments and their interaction with the environment. B. We hypothesize that for many tasks, the global task-level behavior may be decomposed into hierarchically arranged control modules, whose coordinated interactions characterize the dynamics of appendages and body segments. The control module state spaces are embedded as submanifolds (postures) that carry the constituent dynamics (templates) of the full, coordinated system. C. We can use VERT to learn a spatiotemporal filter on the trajectory data that isolates locally stable (template) behavior from transient behaviors reflecting disturbances and stabilizing control (anchoring). D. VERT estimates the distance from each sample to the attractor guiding its asymptotic behavior ($f(z)$, Def. 1). The estimation procedure for a point $z^{*}$ begins with identification of its metric nearest neighbors in state space. E. A Cartesian basis for the tangent space at $z^{*}$ of the manifold containing the trajectories is estimated. The local subsample of the vector field (temporal differences, gray arrows) is projected onto the $n$-dimensional local vielbein basis (black arrows). F. Using this representation, we can estimate the component of the vector field on the vertical subspace, the local trivialization of the global, typically nonlinear vertical subbundle of the attractor. We can use this estimate to approximate the distance from $z^{*}$ to its base point on the unknown attractor $\mathcal{A}$. Steps D-F are repeated for an appropriate number of samples. Filtering the trajectory data for sublevel sets of $f(z)$ reveals the posture carrying the template dynamics (blue) and hierarchical stabilization (anchoring) intervals of what subsequent hypothesis-informed analysis can posit as the stance leg and upper body posture during stance (See Fig. 5).
We consider the setting in which one has collected a dataset of trajectories from a continuous dynamical system. A common hypothesis is that the trajectories of nominally high dimensional and complex dynamical systems lie on or near a lower dimensional, often nonlinear space of latent state variables in the full order observation space, reflecting the presence of, e.g., internal constraints, symmetries, and feedback control. In the context of dynamical systems with multiple spatiotemporal scales, the low dimensional latent submanifolds are typically assumed to be invariant submanifolds of the flow [33]. Using observations to discover latent invariant submanifolds and characterize their dynamics is a central goal of data-driven modeling. By capturing the global structure of the state space, learning algorithms that explicitly infer underlying invariant sets are shown to have superior predictive accuracy, even generalizing to situations where the system is subject to external forces not present in the training data [24].
In principle, one could infer the invariant submanifolds from data by studying the spatiotemporal statistics of the corpus of measured samples. However, this approach implicitly assumes that the samples are in fact drawn (with noise) from the submanifold. While this assumption is justified when we know a priori that the measurements came from steady-state behavior, it does not hold in general for many of the energetically open, nonlinear dynamical systems encountered in nature. Specifically, the presence of transient phenomena resulting from, e.g., external disturbances, noisy internal feedback control processes, and discontinuous events such as the making and breaking of physical contacts will perturb the state away from the latent submanifold and introduce transient dynamics distinct from the on-manifold behavior we want to model.
Because we are interested in experimental observations of physically relevant latent submanifolds, we assume robustness to perturbations in the form of structural stability. The global topology of structurally stable vector fields is invariant to perturbations of the vector field [51], implying that the attracting sets are robust and thus observable even in the presence of noise. Normal hyperbolicity is a sufficient condition for local structural stability; for modeling purposes, we thus assume that detectable attracting sets are normally hyperbolic invariant manifolds (NHIMs) [26]. NHIMs are important to many fields of science and engineering [58], and are central to the modeling of biological locomotion and synthesis of biologically inspired robot controllers [50]. Further discussion of the use of hierarchically composed invariant manifolds in robotics and connections to biological motor control is included in Section 1 of the supplement.
Attracting Submanifolds and Dynamic Locomotion
@@
To realize concrete potential applications of our approach, we consider models of multiscale postural dynamics from dynamic legged locomotion [16]. A longstanding question in the sensorimotor control of locomotion is how the abundance of muscles, joints, and segments of the body are coordinated to achieve a large and diverse set of typically lower dimensional whole body motions. The pioneering work of Nikolai Bernstein [3] suggests that whole body movements are constructed generatively from modular ’submovements’. Modern electromyography studies in both humans and non-human animals support the hypothesis that the neuromuscular system may be organized into modular units (control modules) that can be recruited in series and parallel to achieve high level motor tasks [53]. Development of a systems-level understanding of how such control modules enable robust whole-body behavior is a grand challenge of organismal biology [48], with further applications to robotics and rehabilitation science.
We consider models of multiscale closed loop dynamical systems originally developed at the interface of organismal biology and legged robotics. In this domain, the pairing of a minimal low dimensional latent submanifold and its restriction dynamics with the higher dimensional dynamical system in which it is embedded is called a ’template-anchor pair’ [21]. A typical representation of the (high dimensional) anchor is the full postural state space of an articulated body, consisting of all joints, segments, and actuators. The corresponding (low dimensional) template consists of latent state variables whose spatiotemporal structure and behavior fully characterize the task-level dynamics, e.g., the motion of the center of mass (CoM). A canonical example is that the whole-body dynamics of running are often dynamically consistent with a simple spring-loaded inverted pendulum (SLIP) dynamical system. Here the latent state variables are the generalized coordinates for the effective length of the stance leg and its orientation relative to the ground. Experiments have shown that these latent state variables are sufficient to predict the trajectory of the CoM in animals across a range of and body sizes and morphologies [4], and can be used constructively to stabilize dynamic running behaviors in robots with 2, 4, and 6 legs [29].
Hierarchical Composition of Multiscale Dynamics
To aid in systematic analysis of hierarchical dynamical systems and their state space structures, we refine our notion of hierarchy to the (not mutually exclusive) classes of parallel and sequential dynamical hierarchies from control theory. Compositions of simple dynamical primitives such as point attractors and stable limit cycles have been used extensively as constructive models of both biological and robotic [31, 25, 46, 28] motor control. There is a long history of synthesizing complex whole-body behaviors from hierarchical composition of low dimensional target dynamics, originating with the control design of the first robot capable of dynamic hopping [44]. Parallel hierarchies generalize the template-anchor pairing of the previous section by allowing the embedding and asymptotic stabilization of multiple nested submanifolds in the full order state space. Controllers based on parallel hierarchies of low dimensional dynamics inspired by animal locomotion have been used to stabilize diverse whole body behaviors [12, 13, 54, 22]. More recently, parallel hierarchies have served as an effective inductive bias for learning-based controllers [19, 30]. Sequential hierarchies characterize near instantaneous switches in the governing dynamics as the state is propagated in time. A typical example from biology is the transition between the stance and flight phases of running. A common instance of sequential hierarchy in robotics is the back-chaining of basins of attraction around successively targeted attracting sets to achieve goal states while avoiding obstacles, either as a design choice [37] or as a consequence of limited control affordance [34].
@@
We introduce the Vielbein Recovery of Templates (VERT) framework, an ILG approach for identifying hierarchically embedded attracting sets from trajectory data. VERT takes a set of observed trajectories and returns the fiberwise proximity of each point to the normally attracting invariant manifold (NAIM) that governs its asymptotic behavior. To establish a theoretical foundation for the computational heuristics used by VERT, we model the trajectories $\{z_{i}(t),t\in[0,T_{i}]\}^{N}_{i=1}$ as samples from the (possibly noisy) flow of a dynamical system $(\mathcal{M},\xi)$, where $\mathcal{M}$ is a piecewise smooth Riemannian manifold of dimension $n$, and $v:\mathcal{M}\rightarrow T\mathcal{M}$ is a piecewise $C^{r\geq 1}$ vector field. We assume the trajectories are embedded in a Euclidean observation space $\mathbb{R}^{D}$.
@@
Let $\mathcal{A}\subset\mathcal{M}$ be an invariant submanifold of dimension $a$ for $(\mathcal{M},v)$: $\phi_{t}(\mathcal{A})\subset\mathcal{A},t>0$, for $\phi_{\cdot}$ being the flow associated to the vector field $v$. We will denote by $U\supset\mathcal{A}$ a tubular neighborhood of this submanifold [27] which is attracted to $\mathcal{A}$. Normal Hyperbolicity implies that the trajectories starting in the neighborhood $U$ converge to $\mathcal{A}$, exponentially fast, uniformly in some fixed Riemannian structure. We can introduce, locally, a privileged coordinate system $z=(x,y)$ on $U$, such that the the coordinate $x$ parameterizes the points of $\mathcal{A}$, while the attractor is given by $\mathcal{A}=\{y=0\}$. The fibers $\mathcal{F}_{x}$ of the projection $(x,y)\mapsto x$ are the analogues of the stable manifolds for the hyperbolic dynamics, trivialized through the selected coordinate system. We notice that this coordinate system does not have to (and in fact, almost never does) extend to a global one. However, its usage provides a convenient modeling setting, and leads to the heuristic assumptions outlined below  [^3] We use the German term Vielbein for frame, a local coordinate system to avoid contamination with the video processing context: the term is used in this meaning in the physics literature..
@@
[…]